# Near-Field Channel Estimation in Dual-Band XL-MIMO with Side Information-Assisted Compressed Sensing

### v1

@@ IV Simulation Results
In our simulation, the dual-band XL-MIMO communication system with $K=32$ subcarriers is considered, while the number of antennas at the BS side is set to $N=400$ in order to meet the near-field requirements in the XL-MIMO systems. For the generation of the out-of-band channel matrix $\underline{\mathbf{H}}$, each entry in $\underline{\mathbf{H}}$ can be obtained by multiplying the corresponding entry in $\mathbf{H}$ by a specific coefficient $Q$, whose amplitude and phase depend on the relationship of the central frequencies between the Sub-6GHz band and the mmWave band, which is denoted as $f_{m}$ and $f_{s}$, respectively [15].
@@
where $\underline{x}_{max}$ and $\underline{x}_{min}$ represent the maximum and minimum in $\{\underline{x}_{1},\underline{x}_{2},\cdots,\underline{x}_{N/d}\}$, respectively, and (19) can be leveraged to generate $\mathbf{p}$ in both Case I and Case II as illustrated in Section II-A. Meanwhile, we consider the more general Case II scenario, where both Sub-6GHz and mmWave bands are within the near-field region, leading to stronger block structure for both channels. The far-field channel, i.e., the Sub-6GHz system in Case I, is regarded as a special case by setting the block length as 1 or a smaller value compared with the near-field scenario. According to (19), $p_{i}$ of the row-block submatrix with the maximum Frobienus norm in $\underline{\mathbf{X}}$ is normalized to 1 in $\mathbf{p}$, and therefore has the value of infinity under the influence of the prior factor $v(p_{i})=D\log(\frac{p_{i}}{1-p_{i}})$. As a result, the CSLW-BOMP algorithm always chooses the row-block submatrix in $\mathbf{X}$ of the mmWave band which corresponds to the strongest row-block submatrix in the support channel $\underline{\mathbf{X}}$ of the Sub-6GHz system, as the starting support information in the first iteration for recovering $\mathbf{X}$, leading to potential performance enhancement of the algorithm. Similarly, the row-block submatrix with the minimum Frobienus norm in $\underline{\mathbf{X}}$ becomes negative infinity due to the existence of $v(p_{i})$, which improves the overall performance of our proposed algorithm by consistently ignoring this worst-performing submatrix. The detailed simulation parameters are listed in Table I, and the conventional OMP and BOMP algorithms [24] are adopted for comparison.
@@
Support band frequency $f_{s}$ 3.5GHz
Support bandwidth $BW_{s}$ 150MHz
Object band frequency $f_{m}$ 28GHz
Object bandwidth $BW_{o}$ 850MHz
@@
In Fig. 3, the channel estimation performance in on-grid scenarios is compared, with the number of row-block matrices $\frac{S}{d}$, number of pilot transmission $M$ fixed to 10 and 50 respectively. The normalized mean-square error (NMSE) is leveraged as the criterion, which can be expressed as
@@
where $\mathbf{X}$ and $\hat{\mathbf{X}}$ represent the angular-domain channel matrix and the estimated angular channel, respectively. Moreover, the optimal bound where the channel-tap selections are set to be correct in each iteration is adopted as the upper bound in on-grid scenarios.
It can be observed that our proposed CSLW-OMP and CSLW-BOMP algorithms outperform the traditional OMP and BOMP algorithms, while the performance of CSLW-BOMP algorithm is close to that of the optimal bound benchmark, indicating the effectiveness of the application of out-of-band spatial information. Moreover, the NMSE performance of BOMP and CSLW-BOMP is also superior to that of the OMP and CSLW-OMP algorithms, since the BOMP and the CSLW-BOMP algorithms take the structural characteristic information into consideration and fully utilizes the block structure, which efficiently enhances the robustness of the algorithm, preventing severe nonzero channel-tap selection errors for assured estimation performance. In addition, the exploitation of side information leads to the continuous decrease of the NMSE of the CSLW-BOMP algorithm as the SNR increases, while other algorithms compared tend to reach a plateau at high SNR scenarios.
@@
As illustrated in Fig. 4, when the channel matrix $\mathbf{H}$ changes from on-grid to off-grid, the NMSE performance varies drastically. Although the channel grid cannot align with the block structure in off-grid scenarios, leading to inherent defect in handling off-grid block structure, the CSLW-BOMP still performs the best between the two block structure-based algorithms due to the assistance of out-of-band spatial information. Meanwhile, the CSLW-OMP algorithm enjoys more desirable performance than other algorithms, since the OMP algorithms can effectively adapt to the off-grid channel features by taking the off-grid structure where $d>1$ as on-grid features where $d=1$, and the CSLW-OMP algorithm can exploit the sophisticated out-of-band spatial information from the Sub-6GHz band to support the mmWave channel estimation.
@@
In Fig. 6, the probability of accurate estimation is simulated as a function of the compression rate $\frac{M}{N}$, which is proportional to the number of required pilot transmissions $M$. It unveils that the probability of accurate estimation increases with the increase of compression rate, and the CSLW-BOMP still enjoys the highest probability of accurate estimation under the same compression rate among all compared algorithms, which is similar to the conclusions in Fig. 5. Meanwhile, when the probability of accurate estimation is fixed, CSLW-BOMP requires the least pilot transmissions, followed by BOMP, CSLW-OMP and OMP. For $100\%$ probability of accurate estimation, CSLW-BOMP can reduce the pilot overhead by 25% when compared to the BOMP algorithm, and 30.1% when compared to the OMP algorithm, substantiating the feasibility of the proposed CSLW-BOMP in low-overhead XL-MIMO communications.
@@ Appendix B Proof of Theorem 2
For the case where $\Delta v<0$, (63) can be obtained by similar deductions. This completes the whole proof. ∎

### v3

@@ IV Simulation Results
In our simulation, the dual-band XL-MIMO communication system with $K=32$ subcarriers is considered. Following the antenna settings in [35, 36, 37, 38, 39], the number of antennas at the BS side is set to $N=256$ on the $f=28$GHz mmWave band, and $\underline{N}=32$ on the $\underline{f}=3.5$GHz Sub-6GHz band. Quantitatively, the antenna aperture can then be calculated as $\frac{Nc}{2f}=1.37$m, which is acceptable in practical use. For the generation of the out-of-band channel matrix $\underline{\mathbf{H}}$, each entry in $\underline{\mathbf{H}}$ can be obtained by multiplying the corresponding entry in $\mathbf{H}$ by a specific coefficient $Q$, whose amplitude and phase depend on the relationship of the central frequencies between the mmWave band and the Sub-6GHz band, which is denoted as $f_{m}$ and $f_{s}$, respectively [15].
@@
where $\underline{x}_{max}$ and $\underline{x}_{min}$ represent the maximum and minimum in $\{\underline{x}_{1},\underline{x}_{2},\cdots,\underline{x}_{N/d}\}$, respectively, and (19) can be leveraged to generate $\mathbf{p}$ in both Case I and Case II as illustrated in Section II-A. Meanwhile, we consider the more general Case I scenario, where both mmWave and Sub-6GHz bands are within the near-field region, leading to stronger block structure for both channels. The far-field channel, i.e., the Sub-6GHz system in Case II, is regarded as a special case by setting the block length as 1 or a smaller value compared with the near-field scenario. According to (19), $p_{i}$ of the row-block submatrix with the maximum Frobenius norm in $\underline{\mathbf{X}}$ is normalized to 1 in $\mathbf{p}$, and therefore has the value of infinity under the influence of the prior factor $v(p_{i})=D\log(\frac{p_{i}}{1-p_{i}})$. As a result, the CSLW-BOMP algorithm always chooses the row-block submatrix in $\mathbf{X}$ of the mmWave band which corresponds to the strongest row-block submatrix in the support channel $\underline{\mathbf{X}}$ of the Sub-6GHz system, as the starting support information in the first iteration for recovering $\mathbf{X}$, leading to potential performance enhancement of the algorithm. Similarly, the row-block submatrix with the minimum Frobenius norm in $\underline{\mathbf{X}}$ becomes negative infinity due to the existence of $v(p_{i})$, which improves the overall performance of our proposed algorithm by consistently ignoring this worst-performing submatrix. The detailed simulation parameters are listed in Table I.
First, numerical results are provided in Fig. 4 to backup analytical results in Theorems 1, 2 and Corollary 1. Since better prior factor $v(p_{i})$ leads to more assured results in minimizing the error probability of wrongly choosing a nonzero channel-tap over a zero one in $\mathbf{X}$, as described in Theorems 1, 2 and Corollary 1, the support recovery accuracy is utilized as the measuring metric to showcase the channel-tap selection performance. Let $M=25$ and $\frac{S}{d}=5$. As shown in Fig. 4, our proposed CSLW-BOMP algorithm outperforms its counterparts, while the performance of the CSLW-OMP algorithm is also better than that of the conventional OMP algorithm, which proves the effectiveness of our derived prior factor $v(p_{i})$ in leveraging the out-of-band spatial information. In addition, with the help of the out-of-band spatial information utilized by $v(p_{i})$, the CSLW-BOMP algorithm is able to estimate the support of the mmWave channel matrix $\mathbf{X}$ more accurately when the SNR is low. At high SNRs, the support recovery accuracy of CSLW-BOMP and conventional BOMP algorithms tends to be the same, since the coefficient $D$ reaches a small value and the CSLW-BOMP slowly converges to the BOMP algorithm as discussed in Theorem 2.
As for the benchmarks in the performance comparison, we adopt various channel estimation methods based on greedy algorithms, beam training and low-rank matrix completion (LRMC) theory, which are listed below. 1) OMP and BOMP: Conventional greedy CS algorithms without any side information [24]. 2) Weighted beam recovery (WBR): The beam selection-based method in [15] using only out-of-band spatial information. 3) Probability-based beam selection (PBS): The beam training based channel estimation algorithm in [40]. 4) Fast alternative least squares (FALS): The two-stage channel estimation algorithm based on low-rank matrix completion theory in [41]. 5) Least squares (LS): Conventional least squares channel estimation, where $\hat{\mathbf{X}}=\mathbf{\Psi}^{\dagger}\mathbf{Y}$.
@@
Support bandwidth $BW$ 150MHz
Object bandwidth $\underline{BW}$ 850MHz
@@
Figure 4: Support Recovery Accuracy versus SNR.
In Fig. 5, the channel estimation performance in on-grid scenarios is compared, with the number of row-block matrices $\frac{S}{d}$, number of pilot transmission $M$ fixed to 5 and 25 respectively. The normalized mean-square error (NMSE) is leveraged as the criterion, which can be expressed as
@@
where $\mathbf{X}$ and $\hat{\mathbf{X}}$ represent the angular-domain channel matrix and the estimated angular-domain channel, respectively. Moreover, the optimal bound where the channel-tap selections are set to be correct in each iteration is adopted as the upper bound in on-grid scenarios. It can be observed that our proposed CSLW-BOMP algorithm outperforms all the other benchmarks, whose NMSE performance is close to that of the optimal bound benchmark, while the CSLW-OMP algorithm also has an advantage in comparison to its conventional counterparts, indicating the effectiveness of the application of out-of-band spatial information. Moreover, the NMSE performance of BOMP and CSLW-BOMP is also superior to that of the OMP and CSLW-OMP algorithms, since the BOMP and the CSLW-BOMP algorithms take the structural characteristic information into consideration and fully utilize the block structure in $\mathbf{X}$, which efficiently enhances the robustness of the algorithm, preventing severe nonzero channel-tap selection errors for assured estimation performance. In addition, the exploitation of both types of side information leads to the continuous decrease of the NMSE of the CSLW-BOMP algorithm as the SNR increases, while other algorithms compared tend to reach a plateau at high SNR scenarios.
@@
As illustrated in Fig. 6, when the channel matrix $\mathbf{H}$ changes from on-grid to off-grid, the NMSE performance of various algorithms varies drastically. Although the channel grid cannot align with the block structure in off-grid scenarios, leading to inherent defect in handling off-grid block structure, the CSLW-BOMP still performs the best between the two block structure-based algorithms due to the assistance of out-of-band spatial information. Meanwhile, the CSLW-OMP algorithm outperforms most of the benchmarks except the FALS method. This is because the OMP algorithms can effectively adapt to the off-grid channel features by treating the off-grid structure, where $d>1$ as on-grid features with $d=1$, as elucidated in Remark 4. Therefore, the CSLW-OMP algorithm can facilitate accurate estimation of the mmWave channel by exploiting the sophisticated out-of-band spatial information from the Sub-6GHz band. Regarding the two-stage FALS benchmark, in order to recover the whole $\mathbf{X}$, certain entries in the channel matrix $\mathbf{X}$ need to be measured as prior knowledge. Since $\mathbf{X}$ is typically not a low-rank matrix, the amount of the required pre-measurements is enormous, leading to unacceptable measurement cost. Numerically, in our settings, at least 65% of the entries in $\mathbf{X}$ need to be measured in the first stage to perform reliable LRMC in the second stage, making the FALS method impractical although it achieves a better performance of about 3dB compared to our proposed CSLW-OMP algorithm.
@@
In Fig. 8, the probability of accurate estimation is simulated as a function of the compression rate $\frac{M}{N}$, which is proportional to the number of required pilot transmissions $M$. It unveils that the probability of accurate estimation increases with the increase of compression rate, and the CSLW-BOMP still enjoys the highest probability of accurate estimation under the same compression rate among all compared algorithms, which is similar to the conclusions in Fig. 7. Meanwhile, when the probability of accurate estimation is fixed, CSLW-BOMP requires the least pilot transmissions, followed by BOMP, CSLW-OMP and OMP. For $100\%$ probability of accurate estimation, CSLW-BOMP can reduce the pilot overhead by 25% when compared to the BOMP algorithm, and $30.1\%$ when compared to the OMP algorithm, substantiating the feasibility of the proposed CSLW-BOMP in low-overhead dual-band XL-MIMO communications.
Figure 9: NMSE versus Amplitude Ratio $C$.
[…]
@@ Appendix B Proof of Theorem 2
For the case where $\Delta v<0$, (63) can be obtained by similar deductions. This completes the whole proof.